# Modeling lower-truncated and right-censored insurance claims with an extension of the MBBEFD class

### v1

@@ 6.5 Power exponentially linked exposure example
Finally, we consider a last example belonging again to the exponentially linked exposure family. This last example is a bit more difficult in handling, but it provides the best results for our example. Choose the power function
@@
This function $b$ is a smooth and strictly concave curve on $[0,1]$. We claim that the condition $b^{\prime\prime}(z)+b^{\prime}(z)^{2}\geq 0$ for all $z\in[0,1]$ is fulfilled, which is necessary in order to obtain a distribution function due to Proposition 5.1. Moreover, the density for $z\in[0,1)$ reads as
@@
The derivative of the density $f_{Z}$ can be obtained using (5.6). It does however not allow to characterize the extrema of the density in this example explicitly. Nevertheless, we observe in Figure 8 that this example contains unimodal densities.
@@
in order to find solutions to the standard and extended maximization problems described in (6.1) and (6.2). Similarly as for the previous examples, we obtain the results shown in Figure 8 and Table 5. This time, the fit seems much better. This is especially the case for the red curve, which represents the solution of the flexible maximization problem. The tail behavior on the right and left sides seem adequate in contrast to the plots presented in the previous examples. This observation is confirmed by the AIC scores attained here, which are lower than the AIC scores of all other examples, i.e., we give clear preference to this last example for this data.
@@ 7 Conclusion
In general insurance, however, we are mainly interested in unimodal skewed densities. Therefore, we extended the MBBEFD class to a much bigger family of distributions that we called the Bernegger class. By starting from the properties of an exposure curve, we introduced two subfamilies, namely the logarithmic and exponentially linked exposure curves. Through various examples, we used the full tractability and flexibility of the Bernegger class in order to fit parameters to general insurance claims using maximum likelihood estimations. It turned out that this large class of distributions contains models allowing to obtain a suitable approximation for the distribution of the used dataset, and in general, we have a rich family of unimodal and skewed densities within the Bernegger class. Going forward, it will be interesting to further characterize and classify the members of the Bernegger class based on different properties. Of course, this might involve exploring other link functions. Another next step is to lift these models to regression models allowing for integrating fixed effects described by covariates.
@@ Appendix A Proofs
(c1) Consider the first case $b<1$. In this case, we have a negative numerator and $b^{1-z}$ is increasing in $z\in[0,1)$. If $0<b^{1-z}<(bg-1)/(g-1)$, we have an increasing derivative, and for $0<(bg-1)/(g-1)<b^{1-z}$, we have a decreasing derivative. Since in this case, $(bg-1)/(g-1)<b<b^{1-z}$ holds for all $z\in[0,1]$, we have a monotonically decreasing density.
(c2) Consider the second case $b>1$. In this case, we have a positive numerator and $b^{1-z}$ is decreasing in $z\in[0,1)$. If $0<b^{1-z}<(bg-1)/(g-1)$, we have a decreasing derivative, and for $0<(bg-1)/(g-1)<b^{1-z}$, we have an increasing derivative. Since in this case, we have $(bg-1)/(g-1)>b>b^{1-z}$ for all $z\in[0,1]$, the density is decreasing.
@@
Proof of Lemma 6.4. Since the function $b$ of this example is strictly decreasing, the necessary condition in order to obtain a well-defined distribution reads $b^{\prime\prime}(z)+b^{\prime}(z)^{2}\geq 0$ for all $z\in[0,1]$ according to Proposition 5.1. This condition holds due to the assumptions made on the different parameters. Indeed, these assumptions imply that $b^{\prime}(z)$ is decreasing, while $b^{\prime\prime}(z)$ is increasing for all $z\in[0,1]$. This means in particular that $b^{\prime\prime}(z)+b^{\prime}(z)^{2}$ is increasing for all $z\in[0,1]$. In order to prove the above necessary condition, it thus suffices to show that
@@
Let us first assume that $\beta>\alpha\epsilon\delta^{\alpha-1}$, which is necessary in order to obtain $b^{\prime}(z)<0$ for all $z\in[0,1]$. This implies that the interior of the absolute value is negative, we can thus write

### v2

@@ 6.5 Power exponentially linked exposure example
We consider a final example belonging to the exponentially linked exposure family. This example is a bit more difficult in handling, but it provides the best results for our dataset. Choose the power function
@@
This function $b$ is smooth and strictly concave on $[0,1]$. We claim that $b^{\prime\prime}(z)+b^{\prime}(z)^{2}\geq 0$ holds for all $z\in[0,1]$, which is a necessary condition in order to obtain a distribution function due to Proposition 5.1. Moreover, the density for $z\in[0,1)$ reads as
@@
The derivative of the density $f_{Z}$ can be obtained using (5.6). It does however not allow to explicitly characterize the extrema of the density in this example. Nevertheless, we observe in Figure 8 that this example contains unimodal densities.
@@
in order to find solutions to the standard and extended maximization problems described in (6.1) and (6.2). Similarly as for the previous examples, we obtain the results shown in Figure 8 and Table 6. This time, the fit seems much better. This is especially the case for the red curve, which represents the solution of the extended maximization problem. The tail behavior on the right and left ends seems adequate in contrast to the plots presented in the previous examples. This observation is confirmed by the AIC scores attained here, which are lower than the AIC scores of all other examples, i.e., we give clear preference to this last example for this data.
@@
6.6 Comparison with the log-normal and gamma distribution
For this real dataset example, we performed an MLE estimation without taking into account any covariates. In that case, performing MLE using lower-truncated and right-censored log-normal and gamma models is feasible. Our aim is to compare the extended fit of the power exponentially linked exposure example to the fit of these classical distributions. For this, we first consider the normalized claim $Z$ and write
$Z=\frac{1}{M}\min\left\{\left(X-d\right)_{+},M\right\}\,|\,X>d=\left.\min\left\{\left(\frac{X}{M}-\frac{d}{M}\right)_{+},1\right\}\,\right|\,\frac{X}{M}>\frac{d}{M},$
where $X$ is assumed to have a log-normal and gamma distribution, respectively. For an arbitrary $M>0$, both distributions share the following scaling property
$X\sim\textrm{LN}(\mu,\sigma^{2})\implies\frac{X}{M}\sim\textrm{LN}(\mu-\log(M),\sigma^{2})\quad\textrm{and}\quad X\sim\Gamma(\gamma,c)\implies\frac{X}{M}\sim\Gamma(\gamma,cM),$
where $\mu\in{\mathbb{R}}$ and $\sigma^{2}>0$ are the parameters of a log-normal distribution, and $\gamma>0$ and $c>0$ are the shape and scale parameters of a gamma distribution. Let us denote by $\theta$ the parameters of the log-normal, respectively, gamma distribution. Due to the above scaling property and by setting $\tilde{d}=d/M$, we can maximize without loss of generality the log-likelihood function
$(\theta,\tilde{d})\mapsto\ell_{Z}(\theta,\tilde{d})=\log(f^{(\theta,\tilde{d})}(Z))\mathds{1}_{\{Z<1\}}+\log(p^{(\theta,\tilde{d})})\mathds{1}_{\{Z=1\}},$ (6.11)
where the absolutely continuous density $f^{(\theta,\tilde{d})}$ is given by
$f^{(\theta,\tilde{d})}(y)=\frac{f_{X}(\tilde{d}+y)}{1-F_{X}(\tilde{d})},$
for $y\in[0,1)$, $\tilde{d}\in(0,\infty)$, and where the point mass $p^{(\theta,\tilde{d})}$ is given by
$p^{(\theta,\tilde{d})}=\frac{1-F_{X}(\tilde{d}+1)}{1-F_{X}(\tilde{d})},$
see (1.2) and (1.3). For both distributions, the number of parameters is thus equal to three and the results of these MLE fits are given in Figure 9 and Table 7. Compared to the extended MLE of the power exponentially linked example, the fit on the absolutely continuous density seems slightly worse for the gamma model, whereas the fit using the log-normal model seems to be accurate for this dataset when looking at Figure 9. These observations are confirmed in Table 7. On the one hand, the gamma model fails to obtain a suitable value for the point mass, and on the other hand, the values obtained for the point mass and the mean are close to the empirical values for the log-normal model, which also achieves the lowest AIC score. Therefore we give preference to the log-normal model over the other models for this dataset.
Figure 9: Comparison with the gamma and the log-normal distribution: densities of the random variable $Z|_{\{Z<1\}}$.
Point mass Mean $\ell_{Z|_{\{Z<1\}}}$ $\ell_{Z}$ AIC
Empirical density (Blue) 0.034 0.339 - - -
Flexible MLE density of the power
exponentially linked example (Black)
0.034 0.339 34 402 15 731 -31 453
MLE density of the log-normal
distribution (Red)
0.031 0.340 34 640 15 956 -31 907
MLE density of the gamma
distribution (Green)
0.024 0.341 34 236 15 355 -30 704
Table 7: Comparison with the gamma and the log-normal distribution: results.
Thus, it seems that the log-normal distribution has a better performance than our examples from the Bernegger class. However, we would like to point out that for fitting the gamma and the log-normal models, we do not work with the original deductible $d$ and maximal cover $M$ of the product since we optimize over $\tilde{d}=d/M$, see (6.11). This means that we do not assume the log-normal and the gamma distribution to fit the total financial losses, but we allow for arbitrary scaling. We also remark that fitting the gamma and log-normal models under the original deductible $d$ and maximal cover $M$ does not lead to competitive models for this dataset. Of course, it is then unclear how one could derive the distribution of the total financial loss from the distribution of the observed insurance claims. We show in the next section that such a derivation is in general not unique and implies to make possibly wrong assumptions on the distribution of the total financial loss.
7 Changes in the deductible or the maximal cover
The primary objective of modeling lower-truncated and right-censored insurance claims is to fit the size of observed claims, or equivalently, the size of financial losses falling in the interval $(d,d+M)$. However, the insurer might also be interested in understanding how a change in the deductible and/or in the maximal cover influences the expected claim size. Let us denote by $\tilde{d}$ the new deductible and by $\tilde{M}$ the new maximal cover. Two cases may arise.
If the deductible increases or the maximal cover decreases such that $(\tilde{d},\tilde{d}+\tilde{M})\subseteq(d,d+M)$, the insurer can use his current claims observations in order to derive a new model for the new range of interest. However, in the case where $(\tilde{d},\tilde{d}+\tilde{M})\not\subseteq(d,d+M)$, e.g., when the deductible decreases and the maximal cover increases, an extrapolation based on the distribution of the observed total financial losses being in $(d,d+M)$ has to be made in order to obtain a candidate for the distribution on the new range of interest. We discuss in this section how the insurer can evaluate the new expected claim size after a change in the deductible or the maximal cover by treating separately the two above cases.
7.1 Increase in the deductible or decrease in the maximal cover
As in the previous sections, we denote the total financial loss by $X$, whereas the insurance claim is denoted by $Y$, see (1.1). Furthermore, we define the normalized insurance claim by
$Z=\frac{1}{M}\min\{(X-d)_{+},M\}\,|\,Z>d.$
After an increase in the deductible or a decrease in the maximal cover such that the new range of interest satisfies $(\tilde{d},\tilde{d}+\tilde{M})\subseteq(d,d+M)$, the new normalized insurance claim becomes
$\begin{split}\tilde{Z}&=\frac{1}{\tilde{M}}\min\left\{(X-\tilde{d})_{+},\tilde{M}\right\}\,|\,X>\tilde{d}\\ &=\frac{M}{\tilde{M}}\left.\min\left\{\left(Z-\frac{\tilde{d}-d}{M}\right)_{+},\frac{\tilde{M}}{M}\right\}\,\right|\,Z>\frac{\tilde{d}-d}{M}.\end{split}$
That is, the distribution of $\tilde{Z}$ is equal to a scaled lower-truncated and right-censored distribution of $Z$, where the lower-truncation point is given by $\bar{d}=(\tilde{d}-d)/M$ and the right-censoring point is $\bar{M}=\tilde{M}/M$ (after subtracting the lower-truncation). The next result shows that if $Z$ belongs to the Bernegger class, then $\tilde{Z}$ also belongs to the Bernegger class. In other words, the Bernegger class is closed under lower-truncation and right-censoring.
Proposition 7.1.
Let $Z$ be a member of the Bernegger class with exposure curve $G$ and let $0\leq\bar{d},\,\bar{M}\leq 1$ such that $\bar{d}+\bar{M}\leq 1$. Moreover, define the scaled lower-truncated and right-censored random variable
$\tilde{Z}=\frac{1}{\bar{M}}\min\{(Z-\bar{d})_{+},\bar{M}\}\,|\,Z>\bar{d}.$
This random variable belongs again to the Bernegger class and its exposure curve is given by
$\tilde{G}:[0,1]\to[0,1],\;z\mapsto\frac{G(\bar{d}+z\bar{M})}{G(\bar{d}+\bar{M})}.$
We point out that in the case where an increase in the deductible and/or a decrease in the maximal cover leads to a new interval satisfying $(\tilde{d},\tilde{d}+\tilde{M})\not\subseteq(d,d+M)$, this last result does not apply since either the lower-truncation point $\bar{d}=(\tilde{d}-d)/M$ is negative or the sum $\bar{d}+\bar{M}=(\tilde{d}-d+\tilde{M})/M$ exceeds $1$. This situation has then to be handled by performing some extrapolation on the observed part of the total financial loss distribution as in the case where the deductible decreases or the maximal cover increases, see next section.
7.2 Decrease in the deductible or increase in the maximal cover
The goal of this last section is to treat the case where the new deductible and the new maximal cover satisfy $(\tilde{d},\tilde{d}+\tilde{M})\not\subseteq(d,d+M)$. This typically happens when the deductible $d$ decreases or when the maximal cover $M$ increases, but also in some cases when $d$ increases and $M$ decreases, as discussed above. Since the insurer only has full knowledge on the distribution of the total final loss on the interval $(d,d+M)$, he will have to perform an extrapolation of the observed density in order to obtain a candidate for the distribution on the new range of interest.
Figure 10: Two possible extrapolations of the density of a lower-truncated and right-censored exponential random variable with $d=2000$ and $M=5000$.
In Example 5.3, we showed that a scaled lower-truncated and right-censored exponential random variable belongs to the Bernegger class. Using this example, we show in Figure 10 that extrapolating a density on a given interval to a larger interval leads to infinitely many possible candidates. For this, we first plot in blue the density of an exponential random variable $X\sim\textrm{Exp}(\lambda)$ for $\lambda=\log(2)/3000$. Then, we set the deductible to $d=2000$ and the maximal cover $M=5000$. The blue area in Figure 10 corresponds to the observable part of the density of the total financial loss of the insurer. Let us denote by $p_{-}$ the probability of $X$ being smaller than $2000$ and by $p_{+}$ the probability of $X$ being larger than $7000$. We plot in red another possible extrapolation of the observed density such that the density of the total financial loss under this new extrapolation is continuously differentiable and such that the area under the red curve at the left and right ends is equal to $p_{-}$ and $p_{+}$, respectively. Using the observations falling in the interval $(d,d+M)$ as well as the point mass in $d+M$ indicating the proportion exceeding this threshold, the insurer may choose either one of the two above extrapolations and, actually, the number of possible candidates is in general infinite. Therefore, we point out that in the case where $(\tilde{d},\tilde{d}+\tilde{M})\not\subseteq(d,d+M)$, the insurer cannot derive the distribution of the total financial loss on this new range without making additional assumptions on the random variable $X$. Using the two-parameter logistic distribution of Example 4.3, the next example shows a possible extrapolation.
Example 7.2 (Lower-truncated and right-censored two-parameter logistic distribution).
Let us assume that $X$ follows a two-parameter logistic distribution as in Example 4.3. Moreover, let $d\in{\mathbb{R}}$ and $M>0$. By Corollary 4.2, the normalized lower-truncated and right-censored random variable $Z$ in (4.5) has as density
$f_{Z}(z)=\frac{M}{\sigma}\left(1+e^{(d-\mu)/\sigma}\right)\frac{e^{(zM+d-\mu)/\sigma}}{\left(1+e^{(zM+d-\mu)/\sigma}\right)^{2}},$ (7.1)
for $z\in[0,1)$, with a point mass in $1$ equal to
$p=\frac{1+e^{(d-\mu)/\sigma}}{1+e^{(M+d-\mu)/\sigma}}.$
Suppose now that we know the values of the $d$, $M$ and $p$ and that we want to retrieve the unknown distribution of the random variable $X$. This distribution is related to the distribution of $Z$ by
[…]
At this point, we have to make the following (possibly wrong) assumptions. We first assume that the random variable $X$ is absolutely continuous and that the support of its density is the whole real line. Moreover, we assume that
$f_{X}(z)=Cf_{Z}\left(\frac{z-d}{M}\right),$ (7.3)
for $z\in{\mathbb{R}}$ and where $C$ is a normalizing constant such that $f_{X}$ becomes a density on ${\mathbb{R}}$. Note that in this last assumption, $f_{Z}$ is now seen as a function defined on the whole real line. In general, it might not be possible to obtain a well-defined function by extending the domain of $f_{Z}$, which is a priori only defined for $z\in[0,1)$, see (7.1). In this case, however, extending the domain of the density $f_{Z}$ to the whole real line leads to a well-defined function. Note that assumption (7.3) might seem natural in view of (7.2). Our above assumptions give us again the original density in this case, namely,
$f_{X}(z)=\frac{e^{(z-\mu)/\sigma}}{\sigma\left(1+e^{(z-\mu)/\sigma}\right)^{2}},\quad\textrm{for}\,-\infty<z<\infty,\,-\infty<\mu<\infty,\,\sigma>0.$
$\blacksquare$
This example shows how we can start from a member of the Bernegger class in order to obtain a candidate for the original distribution. Under some assumptions, we managed to retrieve the original distribution in this case. If we assumed however that the support of the density $f_{X}$ is the positive real line or that $X$ is not an absolutely continuous random variable but rather has some point masses at selected locations, we would have ended with a wrong candidate for the original distribution. Of course, the point mass in $1$ can help us in order to verify whether the made assumptions are plausible but even with this information, the number of possible candidates for the original distribution is in general infinite. That is, we point out again that it is in general impossible to retrieve the original distribution by performing some extrapolation based on the lower-truncated and right-censored distribution.
@@ 8 Conclusion
In general insurance, however, we are mainly interested in unimodal skewed densities. Therefore, we extended the MBBEFD class to a much bigger family of distributions that we called the Bernegger class. By starting from the properties of an exposure curve, we introduced two subfamilies, namely the logarithmic and exponentially linked exposure families. Through various examples, we used the full tractability and flexibility of the Bernegger class in order to fit parameters to general insurance claims using maximum likelihood estimations. It turned out that this large class of distributions contains models allowing to obtain a suitable approximation for the distribution of the used dataset, and in general, we have a rich family of unimodal and skewed densities within the Bernegger class. This class of distribution allows to model lower-truncated and right-censored random variables without making any assumption on the original unobserved distribution. We showed further that it is in general impossible to obtain a unique extrapolation from the observable lower-truncated and right-censored distribution in order to derive the full distribution of the original random variable. That is, from an actuarial perspective, it is not possible to know the distribution of the total financial loss from the insurance claims observations.
Going forward, it will be interesting to further characterize and classify the members of the Bernegger class based on different properties. Of course, this might involve exploring other link functions. Another next step is to lift these models to regression models allowing for integrating fixed effects described by covariates. In this last setup, we point out that relying on numerical approximation for the distribution function in order to perform MLE is not possible for a large number of covariates.
Acknowledgments. We kindly thank the referees for their useful remarks that have helped us to improve this manuscript.
@@ Appendix A Proofs
(c1) Consider the first case $b<1$. In this case, we have a negative numerator in (A.2) and $b^{1-z}$ is increasing in $z\in[0,1]$. If $0<b^{1-z}<(bg-1)/(g-1)$, we have a positive derivative, and for $0<(bg-1)/(g-1)<b^{1-z}$, we have a negative derivative. Since in this case, $(bg-1)/(g-1)<b<b^{1-z}$ holds for all $z\in(0,1]$, we have a monotonically decreasing density.
(c2) Consider the second case $b>1$. In this case, we have a positive numerator in (A.2) and $b^{1-z}$ is decreasing in $z\in[0,1]$. If $0<b^{1-z}<(bg-1)/(g-1)$, we have a negative derivative, and for $0<(bg-1)/(g-1)<b^{1-z}$, we have a positive derivative. Since in this case, we have $(bg-1)/(g-1)>b>b^{1-z}$ for all $z\in(0,1]$, the density is decreasing.
@@
$\Box$
Proof of Proposition 5.4. Let $G$ be an exposure curve belonging to the logarithmic linked exposure family and write
$G(z)=\frac{\log(b(z))-\log(b(0))}{\log(b(1))-\log(b(0))},$
for some twice differentiable function $b:[0,1]\to(0,\infty)$ fulfilling condition (4.2) or (4.3). Then, by defining $m=\min\limits_{z\in[0,1]}\log(b(z))-1$ and $\tilde{b}(z)=\log(\log(b(z))-m)$, we obtain for $z\in[0,1]$,
$\begin{split}G(z)&=\frac{\log(b(z))-\log(b(0))}{\log(b(1))-\log(b(0))}\\ &=\frac{\left(\log(b(z))-m\right)-(\log(b(0))-m)}{(\log(b(1))-m)-(\log(b(0))-m)}\\ &=\frac{\exp\left(\tilde{b}(z)\right)-\exp\left(\tilde{b}(0)\right)}{\exp\left(\tilde{b}(1)\right)-\exp\left(\tilde{b}(0)\right)}.\end{split}$
This means that if the function $\tilde{b}$ fulfills condition (5.2) or (5.3), then the exposure curve $G$ also belongs to the exponentially linked exposure family. The latter holds since for $z\in[0,1]$,
$\tilde{b}^{\prime}(z)$ $=$ $\frac{1}{\log(b(z))-m}\,\frac{b^{\prime}(z)}{b(z)},$
$\tilde{b}^{\prime\prime}(z)$ $=$ $\frac{\log(b(z))-m}{\left[b(z)\left(\log(b(z))-m\right)\right]^{2}}\,\left(b^{\prime\prime}(z)b(z)-b^{\prime}(z)^{2}\right)-\frac{b^{\prime}(z)^{2}}{\left[b(z)\left(\log(b(z))-m\right)\right]^{2}},$
which implies
$\tilde{b}^{\prime}(z)^{2}+\tilde{b}^{\prime\prime}(z)=\frac{\log(b(z))-m}{\left[b(z)\left(\log(b(z))-m\right)\right]^{2}}\,\left(b^{\prime\prime}(z)b(z)-b^{\prime}(z)^{2}\right).$
This means that for any $z\in[0,1]$,
$\tilde{b}^{\prime}(z)^{2}+\tilde{b}^{\prime\prime}(z)\leq 0$ $\iff\quad b^{\prime\prime}(z)b(z)-b^{\prime}(z)^{2}\leq 0,$
$\tilde{b}^{\prime}(z)^{2}+\tilde{b}^{\prime\prime}(z)\geq 0$ $\iff\quad b^{\prime\prime}(z)b(z)-b^{\prime}(z)^{2}\geq 0.$
This shows that any exposure curve belonging to the logarithmic linked exposure family belongs to the exponentially linked exposure family but either $b$ or $\tilde{b}$ may take a more complicated form. The other direction follows by using a similar argument.
@@
Proof of Lemma 6.4. Since the function $b$ of this example is strictly decreasing, the necessary condition in order to obtain a well-defined distribution reads $b^{\prime\prime}(z)+b^{\prime}(z)^{2}\geq 0$ for all $z\in[0,1]$ according to Proposition 5.1. This condition holds due to the assumptions made on the different parameters. Indeed, these assumptions imply that $b^{\prime}(z)$ is negative and decreasing, while $b^{\prime\prime}(z)$ is increasing for all $z\in[0,1]$. This means in particular that $b^{\prime\prime}(z)+b^{\prime}(z)^{2}$ is increasing for all $z\in[0,1]$. In order to prove the above necessary condition, it thus suffices to show that
@@
By assumption, we have that $\beta>\alpha\epsilon\delta^{\alpha-1}$. This implies that the interior of the absolute value is negative, we can thus write
@@
Proof of Proposition 7.1. Using (2.1), the distribution of $\tilde{Z}$ can be written as
[…]
for $z\in[0,1]$. By defining a new exposure curve $\tilde{G}:[0,1]\to[0,1],\,z\mapsto G(\bar{d}+z\bar{M})/G(\bar{d}+\bar{M})$ and using Theorem 2.1, we can derive the distribution function induced by $\tilde{G}$, which reads as
$\begin{split}\tilde{F}(z)&=\left(1-\frac{\tilde{G}^{\prime}(z)}{\tilde{G}^{\prime}(0)}\right)\mathds{1}_{\{z\in[0,1)\}}+\mathds{1}_{\{z=1\}}\\ &=\left(1-\frac{G^{\prime}(\bar{d}+z\bar{M})}{G^{\prime}(\bar{d})}\right)\mathds{1}_{\{z\in[0,1)\}}+\mathds{1}_{\{z=1\}}.\end{split}$
This distribution is equal to the distribution of $\tilde{Z}$, which shows that $\tilde{Z}$ is a member of the Bernegger class with exposure curve $\tilde{G}$.
$\Box$
Appendix B MLE parameters of the models of Section 6
We provide in this appendix the values of all MLE parameters obtained when fitting the real dataset of Section 6 in the following tables.
MLE parameters $b$ $g$ $q$
Standard MLE density (Green) $5.66\cdot 10^{-3}$ $50.5$
Flexible MLE density (Red) $2.29\cdot 10^{-3}$ $94.9$ $0.0339$
Table 8: MLE parameters of the MBBEFD example rounded to three significant digits (Section 6.1).
MLE parameters $\alpha$ $\delta$ $a$ $q$
Standard MLE density (Green) $3.76\cdot 10^{4}$ $1.95\cdot 10^{5}$ $0.393$
Flexible MLE density (Red) $9.32\cdot 10^{3}$ $5.67\cdot 10^{4}$ $0.275$ $0.0339$
Table 9: MLE parameters of the power logarithmic linked exposure example rounded to three significant digits (Section 6.2).
MLE parameters $\alpha$ $\beta$ $a$ $q$
Standard MLE density (Green) $2.57$ $-1.25$ $1.05$
Flexible MLE density (Red) $1.94\cdot 10^{-3}$ $-1.57$ $1.00$ $0.0339$
Table 10: MLE parameters of the sine logarithmic linked exposure example rounded to three significant digits (Section 6.3).
MLE parameters $\alpha$ $\beta$ $q$
Standard MLE density (Green) $-2.19$ $-2.88$
Flexible MLE density (Red) $-3.28$ $-3.10$ $0.0339$
Table 11: MLE parameters of the quadratic exponentially linked exposure example rounded to three significant digits (Section 6.4).
MLE parameters $\alpha$ $\beta$ $\delta$ $\epsilon$ $q$
Standard MLE density (Green) $1.00$ $-191$ $0.130$ $-195$
Flexible MLE density (Red) $1.00$ $-3000$ $0.235$ $-3000$ $0.0339$
Table 12: MLE parameters of the power exponentially linked exposure example rounded to three significant digits (Section 6.5).
MLE parameters $\mu$ $\sigma$ $\tilde{d}$
MLE density $-0.947$ $0.563$ $0.110$
Table 13: MLE parameters of the lower-truncated and right-censored log-normal distribution rounded to three significant digits (Section 6.6).
MLE parameters $\gamma$ $c$ $\tilde{d}$
MLE density $1.97$ $5.46$ $0.0154$
Table 14: MLE parameters of the lower-truncated and right-censored gamma distribution rounded to three significant digits (Section 6.6).